Jon wins 5 bets every day for 30 days which equal to 150 bets a month. His dilemma is that every 20 bets he wins, he lost 7 bets. So, how many bets does he needs to win in order to win 150 bets for the month?
Answer #1

He need to win 150/(20-7) * 20 = 230.76 = 231 bets per month in order to win 150 bets.

As for every 20 bet he loses 7, so number of winning bets = 20-7 = 13. Number of rounds of such 20 bets require : 150/13

Each round has 20 bets so total bets = 150*20/13 = 230.76